OpenStudy (anonymous):

The sum of the first 17 terms of an arithmetic sequence is 1,360. If the first term is 16, find the 17th term.

OpenStudy (anonymous):

S_n = n/2 (a_1 + a_n) 1360 = 17/2 (16 + a_17) a_17 = 144.